### v1

# DisasterM3: A Remote Sensing Vision-Language Dataset for Disaster Damage Assessment and Response

Abstract
[…]
1 Introduction
Onset natural and man-made disasters represent one of humanity’s greatest challenges, causing devastating impacts across national borders [46, 8]. These catastrophic events (including earthquakes, tsunamis, floods, explosions, storms, etc) claim tens of thousands of lives globally each year while causing massive infrastructure damage and economic losses [33, 25]. Remote sensing (RS), as an ultra-long-distance Earth observation technology, has been widely used in disaster scenarios, i.e, hurricane damage assessment [29], landslide detection [36], mapping of burn area and ecological impacts [26], etc. Considering the urgency and timeliness of disaster relief, developing AI-based algorithms is necessary.
The recent advent of large vision-language models (VLMs) [16, 42, 6] has achieved substantial milestones in computer vision due to their exceptional ability to reason about visual and linguistic clues and summarize high-level human-readable text. Inspired by the success of the generic domain, remote sensing has also explored the applications of VLMs, i.e., image classification [14], image captioning [13], visual question answering [41], etc. These remote sensing-tailored VLMs show great potential as general-purpose task solvers for multi-task scenarios. Unlike existing research that primarily addresses general geospatial tasks, our work explores the reasoning capabilities of VLMs in extreme disaster scenarios, thereby supporting rescue teams and planning personnel in making informed decisions.
Figure 1: Task taxonomy in DisasterM3 dataset. Each scene includes the paired pre- and post-disaster images. The modalities of post-disaster images are optical or SAR. The 9 tasks derive from 5 essential capabilities for bi-temporal disaster assessment and response: recognition, counting, localization, reasoning, and report generation.
To facilitate the development of VLMs in disaster response, we propose the DisasterM3 dataset, featuring multi-hazard, multi-sensor, and multi-task challenges. As shown in Fig. 1, the DisasterM3 dataset includes the co-registered pre- and post-disaster optical and SAR images, as well as disaster instruction pairs. Our motivation is that a well-performing VLM should possess the ability to achieve a comprehensive understanding of disaster scenarios by responding to the instructions of rescuers. Based on this assumption, we build our task taxonomy by summarizing five essential capabilities required for disaster assessment and response: disaster recognition, damage counting, reasoning and localization, and disaster report generation. Then, these capabilities are delineated with 9 disaster-related tasks, carefully aligning with assessment and response requirements. The diversity of scenarios is ensured by meticulously collecting images from 36 disaster events covering 5 continents. A comprehensive data cleaning, annotation repurposing, instruction design, manual verification, and sampling pipeline is leveraged to generate high-quality annotations. After extensive benchmarking experiments, we found that the cutting-edge VLMs struggle on disaster tasks. Four VLMs were fine-tuned using our dataset and achieve stable improvements across different tasks and sensors, providing solid baselines. The main contributions of this paper could be summarized as follows:
  1. 1.
    To advance intelligent disaster response, we introduce DisasterM3, a multi-hazard, multi-sensor, and multi-task remote sensing dataset for vision-language understanding. It includes 26,988 bi-temporal optical and SAR images, 123,010 instruction pairs across disaster bearing-body recognition, structural damage assessment, referring segmentation, object relational reasoning, comprehensive disaster description, and restoration advice generation.
  2. 2.
    To systematically analyze the efficacy of existing models on disaster tasks, we benchmark 14 advanced large VLMs, including open-source, commercial, and remote sensing methods. The comparative and detailed analysis illuminates their capabilities while identifying several critical directions for future improvement in disaster-focused vision-language understanding.
  3. 3.
    To provide strong baselines, we fine-tune Qwen2.5-VL, InternVL3, LISA, and PSALM on the DisasterM3 dataset, achieving consistent performance enhancements across all evaluation tasks and sensor modalities. With the injection of disaster corpus, the fine-tuned models exhibit good stability to prompt variations, serving as solid baseline solutions.
2 Related Work
General Vision-language Model. Assisted by the strong reasoning abilities of large language models, VLMs have transformed the visual perception domain by enabling the interpretation and reasoning about images through natural language interfaces. Several leading VLMs, including Flamingo [1], MiniGPT-4 [51], LLaVA [17], LLaVA-OneVision [16], InstructBLIP [6], and Qwen2-VL [42], have achieved remarkable results on vision-language tasks. However, these models are limited to generating only textual outputs that describe the image holistically. This restricts their applicability in damage assessment tasks that require the pixel-level detailed understanding. Several approaches have emerged to extend VLMs with fine-grained visual understanding. Ferret [47], Kosmos-2 [27], and VisionLLM [43] incorporate grounding functionalities through bounding box coordinate regression. Besides, LISA [15], PixelLM [31], GLaMM [30], and PerceptionGPT [28], integrate mask decoders to generate object masks from specialized tokens. For richer representation, PSALM [50] and HyperSeg [44] leverage queries in Mask2Former for unified segmentation. Despite their capabilities, generic VLMs exhibit substantial limitations in disaster scenarios due to insufficient domain-specific knowledge, restricting their operational utility in emergency response applications.
Table 1: Comparison of DisasterM3 with existing remote sensing vision-language datasets.
Dataset Propose #Optical #SAR #MT pairs #Text Recognition Counting Localization Reasoning Caption
RSICD [23] General 10,921 - - 54,605 -
RSICap [11] General 2,585 - - 2,585 -
DIOR-RSVG [49] General 17,402 - - 38,320 Pixel
RRSIS-D [20] General 17,402 - - 17,402 Pixel
RSVQA-HR [22] General 10,659 - - 1,066,316 -
EarthVQA [41] General 6,000 - - 208,593 -
RSIEval [11] General 100 - - 933 -
VRSBench [18] General 29,614 - - 205,317 Box
XLRSBench [39] General 1,400 - - 45,942 Box
GeoChatSet [14] General 106,747 - - 308,861 Box
TeoChatlas [13] General 351,957 - 245,210 554,071 Box
FloodNet [29] Disaster 2,348 - - 7,345 -
DisasterM3 (Ours) Disaster 22,214 4,774 15,881 123,010 Pixel
Remote Sensing Vision-language Dataset. Following the substantial progress of general VLMs, the RS field has likewise undergone accelerated development, accompanied by the emergence of numerous specialized vision-language datasets. Focusing on holistic analysis, EarthVQA [41] and RSIEval [11] datasets provide manual instructions for visual question answering (VQA) and image captioning tasks. Leveraging GPT-4, VRSBench [18] introduced visual grounding tasks to evaluate the object reasoning abilities and XLRSBench [39] focuses on ultra-high-resolution image understanding. GeoChatSet [14] and TeoChatlas [13] collect the existing classification and detection datasets for secondary development, formulating the unified instruction-following datasets. Although TeoChatlas involves some disaster scenes, the instructions focus on common recognition tasks. FloodNet [29] is a VQA disaster dataset that assesses the buildings and roads affected by Hurricane Harvey. Limited by its single disaster and simple tasks, it is difficult to fully unleash the potential of VLMs. Overall, RS visual-language datasets for general geospatial tasks have reached a considerable level of maturity, yet there persists a notable deficiency in datasets addressing specialized geoscience challenges. For this case, we design the DisasterM3 dataset that is tailored for global disaster assessment and response with multi-sensor images, bi-temporal inputs, refined damage masks, and diverse visual understanding tasks in the context of disaster.
3 DisasterM3 Dataset
As shown in Fig. 2, we collect 36 historical natural and man-made significant disasters to construct the DisasterM3 dataset. There are 26 events from the xBD [9] and BRIGHT [5] dataset, we extend 10 new events using Maxar’s Open Data program [24]. Considering these optical sensors (WorldView series) have similar spatial resolutions, all pre- and post-disaster images were pre-processed into 0.8m. We collect the post-disaster Synthetic Aperture Radar (SAR) images from Capella Space [4] and Umbra [37]. Considering the amplitude data in the VV or HH bands, SAR images were terrain-corrected, stretched into [0, 255], and finally resampled to match the optical resolution. We performed the georeferencing to ensure that the pre- and post-disaster image pairs are strictly aligned spatially. Following the United Nations Satellite Centre (UNOSAT) Emergency Mapping Products [38], and the Federal Emergency Management Agency (FEMA) [7], we design 9 essential tasks required for disaster assessment and response, evaluating the VLM performances from different aspects.
Figure 2: The DisasterM3 dataset involves 36 significant natural and man-made disaster events (10 types) across five continents. Diverse disaster-centric tasks provide a comprehensive evaluation benchmark for VLMs.
3.1 Perception and Reasoning Tasks in the Context of Disaster
Disaster Recognition. The disaster recognition tasks provide a brief description of disaster scenes, i.e., disaster types, land-use types and key disaster-bearing bodies. The disaster type follows the official definition and and we chose 13 common land-use types from the AID dataset [45] for annotation. The land-use answers include: airport, bridge, river, forest, low vegetation, pond, parking, port, viaduct, residential area, industrial area, commercial area, and sea. Disaster-bearing bodies are the key resources that are damaged by disasters [8], and we focus on 12 types, i.e., building, stadium, open-space ground, bridge, dam, road, port facility, storage tank, farmland, forest, coastline, and mining area. Based on basic recognition types, users could have a rough disaster profile.
Damage Assessment. The damage assessment provides a quantitative analysis of disaster-bearing bodies. We chose the road and building, two important man-made structures for damage assessment. We annotate instance-level building damage masks using ‘intact’, ‘damaged’, and ‘destroyed’ types following FEMA guidelines. As a critical transportation hub, road accessibility plays a vital role in emergency response and recovery efforts. We classify the damaged roads into three types, i.e., ‘intact’, ‘flooded (blocked by water)’, and ‘debris covered (blocked by debris)’. Based on these damage masks, the building counting and road area estimation instructions were automatically generated. The imbalanced sample distributions of damaged buildings and roads (Fig. 2) reveal the actual challenges for model optimization.
Figure 3: Disaster referring segmentation task involves disaster-bearing body mapping and risk analysis. By querying, rescuers could accurately locate the disaster-related objects.
Disaster Referring Segmentation. Each disaster includes different forming factors and prone environments. In addition to disaster-bearing-body mapping, we identify the key visual objects and perform risk analysis using referring segmentation. As shown in Fig. 3, the first example shows an earthquake scene. In addition to referring segmentation for disaster-bearing bodies, we also design the task for finding the optimal rescue places shown in Fig 3(d). Similarly, Fig 3(e) shows the place where rescuers could find the available vehicles for dispatch. As for the volcano eruption scene, we set the instruction tasks to individually map damaged buildings and roads, as well as the lava. Considering the situation, the intact buildings near the lava are also required for segmentation. By polygon distance analysis using the ArcGIS toolbox, the intact buildings within a 100-meter proximity to lava are segmented, providing early warning information. All the referring segmentation tasks are designed according to the specific disaster scenarios, which enable the rescuers to accurately locate the disaster-related objects and places.
Damaged Object Relational Reasoning. To capture the spatial relationships between multiple damaged objects, relational reasoning tasks are designed. In Fig. 4 wildfire scene, the spatial relationships between unaffected buildings and refuge squares, as well as between burnt grassland and unaffected trees, reveal crucial patterns in disaster response and spread prevention. The war conflict scene depicts the damaged industrial area, where the relationships between key facilities, factories, and transportation hubs are clarified. The reasoning task provides spatial analysis services for multiple objects, helping rescuers to understand critical facility spatial dependencies.
Figure 4: Damaged object relational reasoning task describes spatial relationships between key facilities, revealing crucial patterns in the object dependencies.
Figure 5: (Left) The disaster comprehensive reports provide a holistic analysis of disaster situations and evidence-based rescue support. It is notable that immediate earthquake response prioritizes deploying temporary shelters within the stadium for displaced survivors, an intervention demonstrated in the post-disaster image. (Right) Word cloud of reports shows that the disaster-centric words have a considerable degree of diversity.
Disaster Comprehensive Report. To go beyond traditional perception tasks, the comprehensive reports are designed for the holistic analysis of disaster situations. Fig. 5 shows two samples of disaster caption and recovery advice. The earthquake caption describes the collapsed buildings and blocked roads, causing severe traffic congestion. Immediate response advice prioritizes the deployment of temporary shelters within the stadium for displaced survivors, a recommendation visibly implemented in the post-disaster image. Long-term recovery focuses on earthquake-resistant strategies in rebuilding and disaster protocols to mitigate seismic risks. The flooding caption summarizes that roads, buildings, and natural areas experienced severe inundation, while water bodies expanded and merged with flooded regions. Correspondingly, repairing critical transportation infrastructure, establishing temporary residential facilities, and implementing disease prevention protocols are proposed as immediate response measures. The installation of drainage systems integrated with local hydrological networks is recommended as a long-term strategy. Fig. 5 (e) and (f) shows the word cloud of disaster reports. Thanks to the wide range of disaster types, the words are diverse in terms of both nouns and verbs. Most words are disaster-centric, describing bearing bodies, damage impacts, response strategies, etc. Comprehensive disaster reports equip rescuers with enhanced situational awareness and evidence-based decision support.
3.2 Dataset Construction Pipeline
Following the common vision-language data pipeline [21, 48], we divided the whole dataset into Instruct (17,190 Optical images, 3,798 SAR images, and 92,968 instruction pairs) and Bench sets (5,024 Optical images, 976 SAR images, and 30,042 instruction pairs). We describe the detailed annotation process in Fig. 6. As for recognition tasks, GPT-4o was employed to generate diverse prompt variations with similar semantic intent. Disaster domain experts subsequently annotate correct answers for these prompts. These question-answer pairs constitute the Instruct training set. To formulate the multiple-choice Bench set, correct answers were combined with other options. Regarding counting tasks, we counted semantic polygons using annotated road and building damage masks, generating correct answers. The similar options are generated with controlled deviations (±20% and ±40%) to maintain plausibility. As for relational reasoning, the experts annotate bounding boxes and describe the concrete relationship. We use GPT4-o to analyze the image by listing other significant relationships, generating alternatives. As for disaster reports, by referring to bi-temporal images and all basic task information, multiple experts draft the disaster caption and restoration advice following goals of UNITAR and and FEMA projects. GPT4 then polished the reports and corrected grammar errors. Finally, the multi-round verification was performed for controlling quality (Appendix § A). As for uninterpretable parts of SAR imagery, we annotate answers using co-registered optical images and then apply the instructions to SAR images. Using this pipeline, more future disasters can be effectively extended for DisasterM3 dataset.
Figure 6: Dataset construction pipeline. We conduct a thorough process of question designing, answer annotation, and generation of similar questions as well as other options. Multi-round inspection controls the quality of each construction step.
4 Benchmark Experiments
Implementation Setting. As the DisasterM3 dataset features multi-sensor and multi-task, we comprehensively benchmark VLMs under four settings: Optical-Optical and Optical-SAR QA tasks, as well as Optical-Optical and Optical-SAR referring segmentation tasks.. As for QA tasks, LLaVA-OneVision [16], InternVL3 [53], Kimi-VL [35], and Qwen2.5-VL [3]. In addition, we also tested commercial models such as GPT4o [12] and Claude-3 [2] for comparison with the open-source models. As for remote sensing VLMs GeoChat [14], TeoChat [13], EarthDial [34] are chosen for evaluation. As for referring segmentation models, generic VLMs models such as LISA [15], PSALM [50], and HyperSeg [44], alongside the remote sensing model GeoPixel [32] were benchmarked. We fine-tuned Qwen2.5-VL-7B, InvernVL3-8B, LISA and PSALM on our Instruct set. Model details are provided in Appendix § B.
Evaluation Metrics. Following common settings [16, 34], we adopted accuracy (%) for the multiple-choice tasks, i.e., disaster scene recognition (DSR), disaster type recognition (DTR), bearing body recognition (BBR), damaged building counting (DBC), damaged road estimation (DRE), object relational reasoning (ORR). The open-ended tasks are scored using GPT4.1 at a scale of 5 points. Disaster caption is measured from damage assessment precision (DAP), damage detail recall (DDR), and factual correctness (FC). Restoration advice is measured from recovery necessity (RN), strategic completeness (SC), and action priority precision (APP). The average accuracy (AVG) denotes the overall performance. Evaluation prompts are provided in Appendix § C. As for referring segmentation, we chose cIoU and mIoU following previous work [15, 50].
4.1 Comparative Results
Domain gap for disaster scenarios. Tab. 2 presents performance evaluations on optical-optical settings for QA tasks. As a traditional VLM, LLaVA-1.5 exhibited significant limitations when processing disaster scenes due to the domain gap. By leveraging extensive multi-modal pretraining datasets and implementing the AnyRes architecture, LLaVA-OV demonstrates enhancements in both accuracy and multi-image processing capabilities. As efficient Mixture-of-Experts (MoE) VLMs, Kimmi-VL-Think exceeds Kimmi-VL-Instruct in mathematical counting tasks (BDC, DRE). However, the non-negligible domain gap limits their application on complex tasks, particularly degrading performance to near-random levels on the ORR task. This motivated our development of the DisasterM3 dataset, which identifies performance gaps through the Bench set while providing complementary training data via the Instruct set.
Table 2: Benchmarking results of VLMs on DisasterM3 Bench set with optical-optical setting.
Method Accuracy (%) Disaster Caption Restoration Advice
AVG LUC DTR BBD BDC DRE ORR AVG DAP DDR FC AVG RNR APP SC
Random Guess - - 20 - 20 20 20 - - - - - - - -
[…]
LLaVA-1.5-7B [19] 12.1 4.2 - - - - 20.0 - - - - - - - -
LLaVA-OV-7B [17] 24.5 16.3 53.5 3.7 26.4 24.2 22.7 1.66 1.50 1.53 1.93 2.30 3.01 2.08 1.81
Kimi-VL-A3B-Instruct [35] 25.6 28.9 66.3 4.0 20.4 15.0 18.9 1.69 1.53 1.72 1.81 2.67 3.57 2.40 2.05
Kimi-VL-A3B-Think [35] 26.7 27.0 51.6 7.4 24.4 25.4 24.4 1.61 1.39 1.68 1.75 2.61 3.35 2.34 2.15
InternVL3-8B [53] 31.3 39.6 53.5 4.0 30.3 24.1 36.2 1.96 1.88 1.92 2.09 2.75 3.52 2.53 2.21
InternVL3-14B [53] 35.7 42.5 62.0 4.9 27.4 23.6 54.1 2.08 2.01 2.01 2.22 2.86 3.67 2.62 2.29
InternVL3-78B [53] 39.3 43.5 72.5 5.3 29.4 28.7 56.1 2.79 2.74 2.75 2.89 2.90 3.64 2.64 2.43
Qwen2.5-VL-3B [3] 26.2 30.8 56.1 5.7 29.9 21.2 13.8 1.00 0.83 1.05 1.12 2.15 2.98 1.77 1.71
Qwen2.5-VL-7B [3] 31.2 28.3 66.6 4.7 34.2 29.3 23.9 1.75 1.69 1.71 1.85 1.95 2.53 1.83 1.49
Qwen2.5-VL-32B [3] 35.3 36.7 54.7 11.6 33.2 30.9 44.8 1.55 1.42 1.52 1.72 2.96 3.63 2.71 2.55
Qwen2.5-VL-72B [3] 40.5 47.0 74.8 6.8 34.8 28.9 50.8 2.01 1.99 2.00 2.05 2.92 3.79 2.70 2.27
GeoChat-7B [14] 10.7 6.1 - - - - 15.3 - - - - - - - -
TeoChat-7B [13] 23.0 6.9 64.9 2.0 22.5 23.3 18.2 1.77 1.61 1.74 1.96 1.95 2.59 1.77 1.49
EarthDial-4B [34] 22.9 10.6 58.1 3.2 30.2 20.8 14.5 1.53 1.22 1.64 1.73 2.42 3.21 2.08 1.98
[…]
GPT4o [12] 39.3 49.4 80.5 10.6 24.2 21.4 49.8 2.27 2.25 2.28 2.28 3.19 3.92 2.95 2.69
GPT4.1 [12] 42.3 52.4 79.6 7.2 25.5 25.0 64.0 2.57 2.60 2.58 2.54 3.14 3.94 2.93 2.56
[…]
Qwen2.5-VL-7B [3] 40.4 37.7 83.6 21.5 34.3 29.4 36.2 3.90 3.76 3.53 4.41 3.11 3.73 2.88 2.73
$\Delta$uparrow9.2 $\uparrow$9.4 $\uparrow$17.0 $\uparrow$16.8 $\uparrow$0.1 $\uparrow$0.1 $\uparrow$12.3 $\uparrow$2.15 $\uparrow$2.07 $\uparrow$1.82 $\uparrow$2.56 $\uparrow$1.26 $\uparrow$1.20 $\uparrow$1.83 $\uparrow$1.24
InternVL3-8B [53] 41.7 42.6 79.3 23.9 29.1 24.9 50.6 3.83 3.69 3.49 4.32 3.31 3.92 3.10 2.90
$\Delta$uparrow10.4 $\uparrow$3.0 $\uparrow$25.8 $\uparrow$19.9 $\downarrow$-1.2 $\uparrow$0.8 $\uparrow$14.4 $\uparrow$1.87 $\uparrow$1.81 $\uparrow$1.57 $\uparrow$2.23 $\uparrow$0.56 $\uparrow$0.40 $\uparrow$0.57 $\uparrow$0.69
Larger VLMs achieve higher performances. By scaling up LLMs, InternVL3 and Qwen2.5-VL series demonstrate consistent trends that larger LLMs achieve superior performances, confirming established scaling laws observed in general-domain applications. The commercial models, i.e., GPT4o and GPT4.1, showcase competitive performances across all tasks due to their massive corpus.
Remote sensing VLMs still struggle with disaster tasks. Despite being specifically trained on aerial and satellite imagery, existing remote sensing VLMs exhibit feature representations that inadequately transfer to the unique characteristics of disaster scenarios. DisasterM3 narrows the domain gap by providing specialized disaster-focused vision-language data for Earth observation applications.
Fine-tuned models improve comprehensively. By fine-tuning on DisasterM3 Instruct set, the performances of Qwen2.5-VL and InternVL3 have been significantly improved, narrowing the domain gap. Disaster-specific terminology integration during training significantly enhances report generation quality, resulting in more reasonable and professional reports. However, for building damage counting (BDC) task, the fine-tuned InternVL3 exhibits unexpected performance degradation due to overfitting, and we perform detailed analysis in §4. In the future, object sensitive module [52] and numerical enhanced optimization [41] could be explored for model development.
Figure 7: Benchmarking results of VLMs on DisasterM3 Bench set with optical-SAR setting.
Underrepresentation for SAR images. Disasters are often accompanied by extreme weather, with clouds and rain blocking optical sensors. In this case, the active imaging method SAR can penetrate clouds and fog to obtain accurate surface information. Fig. 7 shows the VLMs’ performances evaluated on paired optical-SAR images. Due to the reduced semantics compared to optical imagery and underrepresentation in generic VLM, the performance using post-SAR images yielded substantially diminished performance across all evaluation tasks. In this scenario, as demonstrated by the Qwen2.5-VL series, increasing LLM size fails to yield stable improvements. The finetuned models alleviate the multi-sensor feature alignment, significantly improving the performances on all tasks. Since there still exists a huge gap compared to the optical-optical setting, more multi-modal pretraining [10] and alignment strategies could be investigated [40] for further improvement.
Table 3: Benchmarking results of referring segmentation VLMs on DisasterM3 Bench set. The accuracies of damage-levels (buildings and roads) are merged for simplification.
Model Optical-Optical (%) Optical-SAR (%)
mIoU cIoU Road Building Other mIoU cIoU Road Building
[…]
PSALM-1.3B [50] 9.7 6.3 2.6 10.2 16.3 8.1 8.8 5.1 11.1
HyperSeg-3B [44] 16.6 14.5 7.5 17.0 25.4 8.8 10.3 4.5 13.1
LISA-7B [15] 27.5 22.1 11.9 25.0 45.6 10.9 12.7 6.0 15.7
GeoPixel-7B [32] 14.3 14.2 8.5 18.1 16.2 4.0 5.1 1.8 6.2
[…]
LISA-7B [15] 44.8 43.7 27.6 41.2 65.5 28.2 29.6 21.5 34.9
$\Delta$uparrow17.3 $\uparrow$21.6 $\uparrow$15.7 $\uparrow$16.2 $\uparrow$19.9 $\uparrow$17.3 $\uparrow$16.9 $\uparrow$15.5 $\uparrow$19.2
PSALM-1.3B [50] 50.5 44.5 30.5 49.1 71.9 31.8 35.2 24.3 39.3
$\Delta$uparrow40.8 $\uparrow$38.2 $\uparrow$27.9 $\uparrow$38.9 $\uparrow$55.6 $\uparrow$23.7 $\uparrow$26.4 $\uparrow$19.2 $\uparrow$28.2
Mask token matters in disaster referring segmentation. Tab. 3 shows the compared results of referring segmentation models with multi-sensor settings. After fine-tuning, LISA and PSALM have achieved significant gains in two settings with the injection of disaster reasoning knowledge during the training. It is notable that PSALM exceeds LISA with much smaller parameters. We attribute this to a more robust mask token representation in PSALM. Unlike LISA, which relies on a single fixed mask token for decoding, PSALM adopts a Mask2Former approach that generates comprehensive mask proposals through multiple mask tokens. We empirically set the number of mask tokens to 100 and observed that performance stabilizes when using more than 50 tokens. Disaster scene referring segmentation usually encompasses diverse objects at varying scales, necessitating robust mask token representations facilitated by LLMs.
4.2 Detailed Analysis
Performance variation across disaster categories. VLM performance exhibits variation across disaster types due to differing disaster causal factors and prone environments. As shown in Tab. 4, all methods demonstrate higher performance on landslide events while achieving notably lower metrics on earthquake, tornado, and explosion scenarios. This is because landslides often occur in rural mountainous regions, presenting simpler scenes with fewer objects. In contrast, the others primarily originate from highly developed urban areas, representing substantially more complex scenes. Due to multi-disaster events, the DisasterM3 dataset could measure VLMs comprehensively with unified metrics for multiple vision-language tasks.
Table 4: Performance variation across disaster categories. Accuracy (%) is calculated for each disaster category across all QA tasks.
Method AVG Landslide Earthquake Tornado Conflict Fire Explosion Tsunami Hurricane Volcano Flooding
LLaVA-OV [17] 21.2 23.6 17.1 19.2 22.8 25.4 18.1 19.5 23.2 23.5 19.9
Kimi-VL-A3B-Instruct [35] 19.9 22.2 17.5 20.1 16.3 26.3 13.2 19.2 21.9 23.5 18.4
Kimi-VL-A3B-Think [35] 22.0 26.4 19.6 21.0 17.4 26.9 16.9 22.6 21.8 25.0 22.3
InternVL3-8B [53] 27.5 41.7 22.2 24.4 21.7 33.0 20.9 28.0 27.3 28.8 27.0
InternVL3-14B [53] 30.0 48.6 22.7 26.6 27.2 33.7 21.1 27.3 29.9 31.3 31.5
InternVL3-78B [53] 31.8 48.6 26.3 27.9 25.0 37.1 25.6 32.0 31.7 32.9 30.9
Qwen2.5-VL-3B [3] 24.5 26.4 19.5 21.9 27.5 32.1 17.9 24.2 24.2 29.6 21.8
Qwen2.5-VL-7B [3] 25.6 34.3 21.0 24.9 17.3 33.7 16.7 28.3 27.8 25.6 25.9
Qwen2.5-VL-32B [3] 31.0 50.0 26.4 27.1 26.6 35.7 23.9 32.8 29.4 30.8 27.7
Qwen2.5-VL-72B [3] 31.8 47.2 25.0 31.1 19.0 39.0 24.0 33.4 34.0 33.9 31.2
GPT4o [12] 30.7 52.8 24.8 25.7 19.6 33.7 25.6 28.0 29.7 35.1 32.0
GPT4.1 [12] 32.4 51.4 26.9 26.7 21.7 35.2 27.7 28.5 33.4 35.8 36.5
[…]
Qwen2.5-VL-7B [3] 32.9 41.7 26.5 30.7 27.7 40.3 22.3 33.0 34.0 41.8 31.1
$\Delta$uparrow7.3 $\uparrow$7.4 $\uparrow$5.5 $\uparrow$5.8 $\uparrow$10.4 $\uparrow$6.6 $\uparrow$5.6 $\uparrow$4.7 $\uparrow$6.2 $\uparrow$16.2 $\uparrow$5.2
InternVL3-8B [53] 34.7 56.9 26.0 31.1 26.1 40.3 27.4 33.1 34.9 39.4 32.2
$\Delta$uparrow7.2 $\uparrow$15.2 $\uparrow$3.8 $\uparrow$6.7 $\uparrow$4.4 $\uparrow$7.3 $\uparrow$6.5 $\uparrow$5.1 $\uparrow$7.6 $\uparrow$10.6 $\uparrow$5.2
Performance biases in VLMs for damage counting. Remote sensing imagery typically encompasses numerous objects exhibiting diverse scales and morphologies, with counting challenges becoming particularly acute when conducting fine-grained damage assessment. Fig. 9 illustrates building damage assessment accuracy as a function of building density within analyzed scenes. For InternVL series models, performance initially declines and then improves as density increases. For peripheral ranges ($<$50 or $\geq$200 buildings), these models demonstrate higher confidence and accuracy. In contrast, GPT-4 models exhibit a clear inverse relationship between building density and accuracy. The fine-tuned InternVL3-8B exhibits substantial improvement in low-density scenes ($<$50 buildings) but notable degradation in other ranges, revealing an overfitting dilemma. Different VLMs have different biases in the damage assessment task. In the future, we can integrate pixel-level semantics provided by the DisasterM3 dataset to alleviate the overfitting risk.
Figure 8: Counting acc v.s. Object density.
Figure 9: Accuracy variation with different prompts.
Effects of different prompts. As shown in Fig. 9, we evaluated the robustness of VLMs with five different prompts, where quartiles, ranges of accuracies are plotted. Due to limited LLM capabilities, LLaVA-OV, TeoChat, EarthDial, and Kimi models exhibit higher sensitivity to prompt variations. Besides, IntvernVL3 and Qwen2.5-VL series models show similar patterns wherein larger LLMs display enhanced stability. In comparison to GPT-4o, GPT-4.1 achieves superior performance with notably improved consistency. Following enrichment with the disaster-specific corpus from the DisasterM3 dataset, th fine-tuned Qwen2.5-VL-7B and InternVL3-8B model demonstrate good stability to prompt variations.
5 Conclusion
Inspired by the rapid development of generic VLMs, the remote sensing vision-language datasets and methods have also been gradually explored. To promote interactive AI disaster response, we propose DisasterM3, a multi-hazard, multi-sensor, and multi-task remote sensing dataset for vision-language understanding. DisasterM3 includes 26,988 bi-temporal images and 123k instruction pairs, 36 disaster events across 5 continents. The comprehensive benchmarking of 14 advanced VLMs elucidates both their capabilities and inherent limitations in disaster contexts. Additionally, through fine-tuning four VLMs with the disaster-specific corpus from DisasterM3, we demonstrate substantial performance enhancements across all evaluation tasks. We believe the proposed dataset and baselines will help bridge the gap in VLM-based disaster applications within Earth vision.
Acknowledgments
This work was supported in part by the Council for Science, Technology and Innovation (CSTI) and the Cross-ministerial Strategic Innovation Promotion Program (SIP) "Development of a Resilient Smart Network System against Natural Disasters" (funding agency: NIED), KAKENHI (25K03145) as well as the NVIDIA Academic Grant Program. This work used computational resources Miyabi supercomputer provided by The University of Tokyo through Joint Usage/Research Center for Interdisciplinary Large-scale Information Infrastructures and High Performance Computing Infrastructure in Japan (Project ID: jh250017). Weihao Xuan is supported by RIKEN Junior Research Associate (JRA) Program. We also thank Ritwik Gupta for sharing the valuable xBD dataset and for his expertise in disaster response guidance.